A very brief introduction to

differentiable manifolds

Tom Carter

http://cogs.csustan.edu/˜ tom/diff-manifolds

Santa Fe Institute Complex Systems Summer School

June, 2001

Our general topics:

Why differentiable manifolds 3 Topological spaces 4 Examples of topological spaces 11 Coordinate systems and manifolds 14 Manifolds 19 References 21

Why differentiable manifolds

• Differentiable manifolds can generally be ← thought of as a generalization of ℝ^{n}. They are mathematical objects equipped with smooth (local) coordinate systems. Much of physics can be thought of as having a natural home in differentiable manifolds. A particularly valuable aspect of differentiable manifolds is that unlike traditional flat (Euclidean) ℝ^{n}, they can have (intrinsic) curvature.

Topological spaces

←

• We need a way to talk about “nearness” of points in a space, and continuity of functions. We can’t (yet) talk about the “distance” between pairs of points or limits of sequences – we will use a more abstract approach. We start with:

Def.: A topological space (X, T) is a set X together with a topology T on X. A topology on a set X is a collection of subsets of X (that is, T ⊂ P(X)) satisfying:

1. If G_{1}, G_{2} ∈ T, then G_{1} ∩ G_{2} ∈ T.

2. If {G_{α} | α ∈ J} is any collection of sets in T, then

_{α∈J}^{⋃} G_{α} ∈ T.

3. ∅ ∈ T, and X ∈ T.

• The sets G ∈ T are called open sets in X. A subset F ⊂ X whose complement is open is called a closed set in X.

• If A is any subset of a topological space X, then the interior of A, denoted by A^{∘}, is the union of all open sets contained in A. The closure of A, denoted by \overline{A}, is the intersection of all closed sets containing A.

• If x ∈ X, then a neighborhood of x is any subset A ⊂ X with x ∈ A^{∘}.

• If (X, T) is a topological space, and A is a subset of X, then the induced or subspace topology T_{A} on A is given by

T_{A} = {G ∩ A | G ∈ T }.

It is easy to check that T_{A} actually is a topology on A. With this topology, A is called a subspace of X.

• Suppose X and Y are topological spaces, and f : X → Y. Recall that if V ⊂ Y, we use the notation

f^{−1}(V) = {x ∈ X | f(x) ∈ V }.

We then have the definition:

Def.: A function f : X → Y is called continuous if f^{−1}(G) is open in X for every open set G in Y.

• We can also define continuity at a point. Suppose f : X → Y, x ∈ X, and y = f(x). We say that f is continuous at x if for every neighborhood V of y, there is a neighborhood U of x with f(U) ⊂ V. We then say that a function f is continuous if it is continuous at every x ∈ X.

• A homeomorphism from a topological space X to a topological space Y is a 1-1, onto, continuous function f : X → Y whose inverse is also continous.

• A topological space is called separable if there is a countable collection of open sets such that every open set in T can be written as a union of members of the countable collection.

• A topological space X is called Hausdorff if for every x, y ∈ X with x ≠ y, there are neighborhoods U and V of x and y (respectively) with U ∩ V = ∅.

• This is just the barest beginnings of Topology, but it should be enough to get us off the ground …

Topological spaces - exercises

1. Show that the intersection of a finite ← number of open sets is open. Give an example to show that the intersection of an infinite number of open sets may not be open.

2. How many distinct topologies are there on a set containing three elements?

3. Show that the interior of a set is open.

Show that the closure of a set is closed. Show that A^{∘} ⊂ A ⊂ \overline{A}. Show that it is possible for A^{∘} to be empty even when A is not empty.

4. Show that if f : X → Y is continuous, and F ⊂ Y is closed, then f^{−1}(F) is closed in X.

5. Show that a set can be both open and closed. Show that a set can be neither open nor closed.

6. Show that if f : X → Y and g : Y → Z are both continuous, then g ∘ f : X → Z is continuous.

7. Show that the two definitions of continuity are equivalent.

8. A subset D ⊂ X is called dense in X if \overline{D} = X. Show that it is possible to have a dense subset D with D^{∘} = ∅.

9. Show that if D is dense in X, then for every open set G ⊂ X, we have G ∩ D ≠ ∅. In particular, every neighborhood of every point in X contains points in D.

10. Show that in a Hausdorff space, every set consisting of a single point x (i.e., {x}) is a closed set.

Examples of topological

spaces

• For any set X, there are two trivial ← topologies:

T_{c} = {∅, X}

and

T_{d} = P(X).

T_{d} is the topology in which each point (considered as a subset) is open (and hence, every subset is open). It is called the discrete topology. T_{c} is sometimes called the concrete topology.

• On ℝ, there is the usual topology. We start with open intervals (a, b) = {x | a < x < b}. An open set is then any set which is a union of open intervals.

• On ℝ^{n}, there is the usual topology. One way to get this is to begin with the open balls with center a and radius r, where a ∈ ℝ^{n} can be any point in ℝ^{n}, and r is any positive real number:

B_{n}(a, r) = {x ∈ ℝ^{n} | |x − a| < r}.

An open set is then any set which is a union of open balls.

Examples of topological spaces

- exercises

←

1. Check that each of the examples actually is a topological space.

2. For k < n, we can consider ℝ^{k} to be a subset of ℝ^{n}. Show that the inherited subspace topology is the same as the usual topology.

3. Show that ℝ^{n} with the usual topology is separable and Hausdorff.

Coordinate systems and

manifolds

←

• Suppose M is a topological space, U is an open subset of M, and μ : U → ℝ^{n}. Suppose further that μ(U) is an open subset of ℝ^{n}, and that μ is a homeomorphism between U and μ(U). We call μ a local coordinate system of dimension n on U.

For each point m ∈ U, we then have that μ(m) = (μ_{1}(m), …, μ_{n}(m)), the coordinates of m with respect to μ.

• Now suppose that we have another open subset V of M, and ν is a local coordinate system on V. We say that μ and ν are C^{∞} compatible if the composite functions μ ∘ ν^{−1} and ν ∘ μ^{−1} are C^{∞} functions on μ(U) ∩ ν(V). Remember that a function on ℝ^{n} is C^{∞} if it is continuous, and all its partial derivatives are also continuous.

• A topological manifold of dimension n is a separable Hausdorff space M such that every point in M is in the domain of a local coordinate system of dimension n. These spaces are sometimes called locally Euclidean spaces.

• A C^{∞} differentiable structure on a topological manifold M is a collection F of local coordinate systems on M such that:

1. The union of the domains of the local coordinate systems is all of M.

2. If μ_{1} and μ_{2} are in F, then μ_{1} and μ_{2} are C^{∞} compatible.

3. F is maximal with respect to 2. That is, if ν is C^{∞} compatible with all μ ∈ F, then ν ∈ F.

• A C^{∞} differentiable manifold of dimension n is a topological manifold M of dimension n, together with a C^{∞} differentiable structure F on M.

Notes:

1. It is possible for a topological manifold to have more than one distinct differentiable structures.

2. In this discussion, we have limited ourselves to C^{∞} differentiable structures. With somewhat more work, we could define C^{k} structures for k < ∞.

3. We have limited the domains of our local coordinate systems to be open subsets of M. This means that the usual spherical and cylindrical coordinate systems on ℝ^{3} do not count as local coordinate systems by our definition.

4. With somewhat more work, we could define differentiable manifolds with boundaries.

5. We have limited ourselves to manifolds of finite dimension. With somewhat more work, we could define infinite dimensional differentiable manifolds.

Coordinate systems and manifolds

- exercises

← 1.

Manifolds ^{←}^{•}

19

Manifolds - exercises

← 1.

←

References

[1] Auslander, Louis, and MacKenzie, Robert E., Introduction to Differentiable Manifolds, Dover Publications, New York, 1977.

[2] Bishop, Richard L., and Goldberg, Samuel I., Tensor Analysis on Manifolds, Dover Publications, New York, 1980.

[3] Warner, Frank W., Foundations of Differentiable Manifolds and Lie Groups, Scott, Foresman and Company, Glenview, Illinois, 1971.

To top ←